How many milliliters of 11.6 M perchloric acid solution should be used to prepare 1.50 L of 0.100 M HClO4?
Answer #1

M1V1 = M2V2

0.100 M *1500 mL = 11.6 M*V2

150 = 11.6 V2

V2 = 150 / 11.6 = 12.93 mL solution is required.